# Essential Dimension of Central Simple Algebras when the Characteristic is Bad

### v1

@@ 6. Sequences of Linked Cyclic Algebras
Here we consider also cyclic algebras of degree $p^{m}$ over $F$ of $\operatorname{char}(F)=p$. Such a cyclic algebra is denoted $\omega,\beta)_{p^{m},F}$ for some $\beta\in F^{\times}$ and $\omega=(\omega_{1},\dots,\omega_{m})\in W_{m}F$ where $W_{m}F$ is the ring of truncated Witt vectors of length $m$ over $F$. The algebra is generated by $u_{1},\dots,u_{m}$ and $y$ subject to the relations $(u_{1}^{p},\dots,u_{m}^{p})-(u_{1},\dots,u_{m})=(\omega_{1},\dots,\omega_{m})$, and $(yu_{1}y^{-1},\dots,yu_{m}y^{-1})=(u_{1},\dots,u_{m})+(1,0,\dots,0)$ where the arithmetic operations follow the rules of Witt vectors and $y^{p^{m}}=\beta$.
It is important to note that these algebras generate ${{}_{p^{m}}Br}(F)$, and that a similar description in terms of differential forms (over $W_{m}F$) exists (see [AJO18]). Certain bounds on the essential dimension extend to the $p^{m}$-torsion setting in a somewhat trivial manner (see for example [CM20, Theorem 5.4]).
@@
The special case of $\operatorname{LCA}_{2,2}$ corresponds to the functor studied in [CR15] of triples of linked quaternion algebras. The essential dimension of $\operatorname{LCA}_{2,2}$ was calculated in [Cha24] to be 3, in accordance with the analogous result from [CR15, Theorem 1.3 (c)]. In [Cha24] it was also shown that $\operatorname{ed}(\operatorname{LCA}_{p,2})=3$.
One key observation is that inseparably linked sequences have small essential dimension. We say that $(C_{1},\dots,C_{n})\in\operatorname{LCA}_{p^{m},n}$ are inseparably linked if they can be written as $C_{1}=[\omega_{1},\beta)_{p^{m},F},\dots,C_{n}=[\omega_{n},\beta)_{p^{m},F}$ for a fixed $\beta\in F^{\times}$. In [CFM23], it was proven that in this case, one can choose $\gamma_{1},\dots,\gamma_{n-1}\in k$ and $\tau\in W_{m}F$ such that $C_{1}=[\tau,\beta)_{p^{m},F},C_{2}=[\tau,\beta+\gamma_{1})_{p^{m},F},\dots,C_{n}=[\tau,\beta+\gamma_{n-1})_{p^{m},F}$.
@@
The first statement is obvious from the result from [CFM23], and the second is immediate from the main result of [RV18]. ∎
@@
In [Cha24] it was shown that $\operatorname{ed}_{p}(C_{1},C_{2})\leqslant 2$ for $(C_{1},C_{2})\in\operatorname{LCA}_{p,2}$ if and only if $C_{1}\otimes L$ and $C_{2}\otimes L$ are inseparably linked for some prime-to-$p$ extension $L/F$.
Question 6.2.
@@
The essential (2-)dimension of $\operatorname{LCA}_{2,n}$ is exactly $n+1$.
@@
By [CL25], there is a well-defined invariant mapping any $n$-tuple $([\alpha,\beta_{1})_{2,F},\dots,[\alpha,\beta_{n})_{2,F})\in\operatorname{LCA}_{2,n}(F)$ to the quadratic $(n+1)$-fold Pfister form $\langle\!\langle\beta_{1},\dots,\beta_{n},\alpha]\!]$. In the generic case, the image is an anisotropic quadratic $(n+1)$-fold Pfister form, so it cannot descend to a field of transcendence degree smaller than $n+1$. Therefore, the essential dimension is no less than $n+1$, and since it is clearly no greater than $n+1$, this is its exact value.
@@
Unfortunately, the symbol $\alpha\frac{d\beta_{1}}{\beta_{1}}\wedge\dots\wedge\frac{d\beta_{n}}{\beta_{n}}$ in $H_{p}^{n+1}(F)$ is not an invariant of the sequence $([\alpha,\beta_{1})_{p,F},\dots,[\alpha,\beta_{n})_{p,F})\in\operatorname{LCA}_{p,n}(F)$ for $p>2$ as shown in [Cha23, Remark 3.6].
@@ 7. The $\operatorname{Dec}_{p^{m},n}$ functor
In this section, we say a few words about the decomposition functor $\operatorname{Dec}_{p^{m},n}$ mapping each field $F$ containing $k$ to the set of isomorphism classes of tensor products of $n$ cyclic algebras of degree $p^{m}$ over $F$. The essential dimension and $p$-dimension of this functor were studied in [McK17] for $m=1$, and it was concluded that they are both equal to $n+1$.
@@
The essential $p$-dimension is at least $n+1$ because of the surjective map from $\operatorname{Dec}_{p^{m},n}(F)$ to $\operatorname{Dec}_{p,n}(F)$ given by mapping each $(\omega_{1},\dots,\omega_{m}),\beta)_{p^{m},F}$ to $\omega_{1},\beta)_{p,F}$. On the other hand, the essential dimension is at most $n+1$ because of the main result of [RV18] that implies that the essential $p$-dimension of $(\mathbb{Z}/p^{m}\mathbb{Z})^{\times n}$ over fields of characteristic $p$ is 1. ∎
@@ 8. Cohomology Groups
In the good characteristic case, Galois Cohomology gives a description of the $n$-torsion part of the Brauer group; ${{}_{n}Br}(F)\cong H^{2}(F,\mu_{n})$. In the bad characteristic case, the analogous cohomological construction is Kato-Milne cohomology for prime-torsion (see [Kat82]) and Izhboldin groups in $p^{m}$-torsion (see [Izh00] and [AJO18]). As pointed out to us by J. P. Tignol, though, they are called “Izhboldin groups”, they were previously known to Kato and used by him (see [Kat80, Section 2, Corollary 4 to Proposition 2]). Both constructions utilize differential forms and are described here.
@@
The construction of Kato-Milne cohomology is as follows: we start with a field $F$ of $\operatorname{char}(F)=p$. Take $\Omega_{F}^{n}$ the ring of $n$-fold differential forms (i.e., sums of terms of the form $a_{1}db_{1}\wedge\dots\wedge db_{n}$ subject to the relations $d(bc)=bdc+cdb$ and $db\wedge db=0$), and consider the map $\wp:\Omega_{F}^{n}\rightarrow\Omega_{F}^{n}/d\Omega_{F}^{n-1}$ defined by $a\frac{db_{1}}{b_{1}}\wedge\dots\wedge\frac{db_{n}}{b_{n}}\mapsto(a^{p}-a)\frac{db_{1}}{b_{1}}\wedge\dots\frac{db_{n}}{b_{n}}$. The group $H_{p}^{n+1}(F)$ is the cokernel of this map, and it is the analogous group to $H^{n+1}(F,\mu_{p}^{\otimes n})$ in the more common case of fields $F$ of characteristic not $p$ containing primitive $p$th roots of unity (see [EKM08, Section 101]). In particular, for $p=2$, $H_{2}^{n+1}(F)$ is isomorphic to $I_{q}^{n+1}(F)/I_{q}^{n+2}(F)$, where $I_{q}^{n+1}(F)$ is the subgroup of the Witt group of nonsingular quadratic forms over $F$ generated by generalized $(n+1)$-fold Pfister forms. The isomorphism is given by $a\frac{db_{1}}{b_{1}}\wedge\dots\wedge\frac{db_{n}}{b_{n}}\mapsto\langle\!\langle b_{1},\dots,b_{n},a]\!]$ (see also [EKM08, Section 16]). For $n=1$, $H_{p}^{2}(F)$ is isomorphic to ${{}_{p}Br}(F)$, and the isomorphism is given by $a\frac{db}{b}\mapsto[a,b)_{p,F}$ (see Section 9). The symbol length of a class in $H_{p}^{n+1}(F)$ is the number of terms of the form $a\frac{db_{1}}{b_{1}}\wedge\dots\wedge\frac{db_{n}}{b_{n}}$ needed to express it, and this number is always bounded from above by $\binom{r}{n}$ when $r$ is the $p$-rank of $F$ (see [CM20, Corollary 3.4]). Knowing the symbol length may say something about the decomposability of the algebra, which is the trick applied in [CM20] to provide lower bounds for the essential dimension of $\operatorname{Alg}_{p^{n},p}$.
@@
The group $\nu(n)_{F}$, defined to be the kernel of $\wp$, is the one corresponding to Milnor $K$-theory (i.e., the group $K_{n}(F)/pK_{n}(F)$), and it is the group $H^{n}(F,\mu_{p}^{\otimes n})$ in the language of [EKM08, Section 101]. For $p=2$, this group is isomorphic to $I^{n}F/I^{n+1}F$ where $I^{n}F$ is the group generated by generalized bilinear $n$-fold Pfister forms inside the additive Witt group $WF$. The Witt group is actually a ring, because the tensor product is defined, and more generally, there is a cup product that maps pairs $(A,B)\in\nu(n)_{F}\times\nu(m)_{F}$ to $A\wedge B\in\nu(n+m)_{F}$. Similarly, there is a cup product that maps pairs $(A,B)\in H_{p}^{n}(F)\times\nu(m)_{F}$ to $A\wedge B\in H_{p}^{n+m}(F)$. But note that since $H_{p}^{n}(F)$ and $\nu(n)_{F}$ need not be isomorphic, there is no cup product for pairs in $H_{p}^{n}(F)\times H_{p}^{m}(F)$. This means that certain easy lower bounds, obtained in the more standard case of characteristic not $p$ for the essential dimension by means of the divided power operators, cannot be applied here. In some sense, the lower bound of $n+1$ obtained in [McK17] for the essential dimension of $\operatorname{Alg}_{p^{n},p}$ is the analogous number to the number one would get in the standard case, because the equivalent lower bound obtained for the generic tensor product of $n$ cyclic algebras of degree $p$ is $2n$ which is obtained by taking the corresponding generic symbol in $H^{2n}(F,\mu_{p}^{\otimes 2n})$.
@@
The Izhboldin groups $H_{p^{m}}^{n+1}(F)$ can be obtained in a similar way as a quotient of $W_{m}\Omega^{n}_{F}$ ([AJO18, Theorem 2.27]), which is the ring of truncated Witt vectors of length $m$ over $\Omega^{n}_{F}$ (see [Jac89, Chapter 8, Section 10]), but a neater description in terms of generators and relations is the following ([Izh00]): take the additive group $W_{m}(F)\otimes\underbrace{F^{\times}\otimes\dots\otimes F^{\times}}_{n\ \text{times}}$ modulo the relations
@@
The group $H_{p^{m}}^{1}(F)$ is of particular interest because it describes the degree $p^{m}$ cyclic extensions of $F$, and is in fact $W_{m}F/\wp(W_{m}F)$. The essential dimension of the functor $H_{p^{m}}^{1}$ is unknown in general, except for the cases of $m=1$ and $2$, for which it is 1 and 2 respectively (see [Led04]). However, the essential $p$-dimension of the functor $H_{p^{m}}^{1}$ is known to be 1 (see [RV18]), which provides for an interesting case where the essential $p$-dimension is strictly smaller than the essential dimension.
@@
Given $(a,b)\in H^{1}_{p^{2}}(F)$, there exist a prime to $p$ extension $L$ of $F$ of degree at most $p^{2}-p+1$ such that $(a,b)_{L}=(c,0)$ for a suitable $c\in L$.
@@
The second slot is thus a monic polynomial of degree $p^{2}-p+1$, and therefore has a root in some prime to $p$ extension $L$ of $F$ of degree at most $p^{2}-p+1$. When $p=2$,
@@
Question 8.2.
Can the argument in Proposition 8.1 be extended to show a similar phenomenon occurs in general for $H_{p^{m}}^{1}$?

### v2

@@ 5. Sequences of Linked Cyclic Algebras
Here we consider cyclic algebras of degree $p^{m}$ over $F$ of $\operatorname{char}(F)=p$. Such a cyclic algebra is denoted $\omega,\beta)_{p^{m},F}$ for some $\beta\in F^{\times}$ and $\omega=(\omega_{1},\dots,\omega_{m})\in W_{m}F$ where $W_{m}F$ is the ring of truncated Witt vectors of length $m$ over $F$.
As a set, $W_{m}F=\{(x_{0},\ldots,x_{m-1})\,|\,x_{i}\in F\}$, but addition and multiplication are defined by Witt Polynomials making them into a ring ([27, Thm.8.30] or [26, pg.22]). For example, $W_{n}=\sum_{i=0}^{n}p^{i}X_{i}^{p^{n-i}}\in\mathbb{Z}[X_{0},\ldots]$ is the $n$-th Witt polynomial and the sum of two truncated Witt vectors $x=(x_{1},\ldots,x_{m})$ and $y=(y_{1},\ldots,y_{m})$ is defined to be $(s_{0}(x,y),s_{1}(x,y),\ldots,s_{m}(x,y))$ where $s_{i}\in\mathbb{Z}[X_{0},\ldots,Y_{0},\ldots]$ is uniquely defined so that $W_{n}(s_{0}(x,y),\ldots)=W_{n}(x)+W_{n}(y)$ for all $n\geqslant 0$. In particular, $s_{0}(x,y)=x_{0}+y_{0}$ and $s_{1}(x,y)=x_{1}+y_{1}+(x_{0}^{p}+y_{0}^{p}-(x_{0}+y_{0})^{p})/p$.
The algebra $\omega,\beta)_{p^{m},F}$ is generated by $u_{1},\dots,u_{m}$ and $y$ subject to the relations $(u_{1}^{p},\dots,u_{m}^{p})-(u_{1},\dots,u_{m})=(\omega_{1},\dots,\omega_{m})$, and $(yu_{1}y^{-1},\dots,yu_{m}y^{-1})=(u_{1},\dots,u_{m})+(1,0,\dots,0)$, where the arithmetic operations follow the rules of Witt vectors and $y^{p^{m}}=\beta$.
It is important to note that these algebras generate ${{}_{p^{m}}{\operatorname{Br}}}(F)$, and that a similar description in terms of differential forms (over $W_{m}F$) exists (see [3]). In [20, Theorem 5.4]) it was shown that $\operatorname{ed}(\operatorname{Alg}_{p^{mn},p^{m}})\geqslant n+1$.
@@
The special case of $\operatorname{LCA}_{2,2}$ provides a characteristic 2 analogue to the functor studied in [13, Theorem 1.3 (c)] of triples of quaternion algebras with trivial tensor product over fields of characteristic 0: each $(Q_{1},Q_{2},Q_{3})$ with trivial $Q_{1}\otimes Q_{2}\otimes Q_{3}$ corresponds to a linked pair $(Q_{1},Q_{2})$ where $Q_{3}$ is Brauer equivalent to $Q_{1}\otimes Q_{2}$. The essential dimension of $\operatorname{LCA}_{2,2}$ was calculated in [16] to be 3, in accordance with the analogous result from [13, Theorem 1.3 (c)]. In [16] it was also shown that $\operatorname{ed}(\operatorname{LCA}_{p,2})=3$.
One key observation is that inseparably linked sequences have small essential dimension. We say that $(C_{1},\dots,C_{n})\in\operatorname{LCA}_{p^{m},n}$ are inseparably linked if they can be written as $C_{1}=[\omega_{1},\beta)_{p^{m},F},\dots,C_{n}=[\omega_{n},\beta)_{p^{m},F}$ for a fixed $\beta\in F^{\times}$. In [18, Theorem 4.7], it was proven that in this case, one can choose $\gamma_{1},\dots,\gamma_{n-1}\in k$ and $\tau\in W_{m}F$ such that $C_{1}=[\tau,\beta)_{p^{m},F},C_{2}=[\tau,\beta+\gamma_{1})_{p^{m},F},\dots,C_{n}=[\tau,\beta+\gamma_{n-1})_{p^{m},F}$.
@@
The first statement follows from the result in [18] mentioned just above the corollary. The second statement follows from the same description of the algebra but using [43, Theorem 1] with the group $G=\mathbb{Z}/p^{m}\mathbb{Z}$ we see that after a prime-to-$p$ extension, the cyclic extension descends to a field of transcendence degree 1. ∎
@@
In [16] it was shown that $\operatorname{ed}_{p}(C_{1},C_{2})\leqslant 2$ for $(C_{1},C_{2})\in\operatorname{LCA}_{p,2}$ if and only if $C_{1}\otimes L$ and $C_{2}\otimes L$ are inseparably linked for some prime-to-$p$ extension $L/F$.
Question 5.2.
@@
Both the essential dimension and the essential 2-dimension of $\operatorname{LCA}_{2,n}$ are equal to $n+1$.
@@
By [19], there is a well-defined invariant mapping any $n$-tuple of linked quaternion algebras $([\alpha,\beta_{1})_{2,F},\dots,[\alpha,\beta_{n})_{2,F})\in\operatorname{LCA}_{2,n}(F)$ to the quadratic $(n+1)$-fold Pfister form $\langle\!\langle\beta_{1},\dots,\beta_{n},\alpha]\!]$. In the generic case, the image is an anisotropic quadratic $(n+1)$-fold Pfister form, so it cannot descend to a field of transcendence degree smaller than $n+1$ over $k$. This is because $k$ is algebraically closed, and every field $E$ of transcendence degree $\leqslant n$ over $k$ has $u$-invariant at most $2^{n}$, and then all quadratic $(n+1)$-fold Pfister forms are isotropic. Therefore, the essential dimension is no less than $n+1$, and since it is clearly no greater than $n+1$, this is its exact value.
@@
The invariant $\langle\!\langle\beta_{1},\dots,\beta_{n},\alpha]\!]$ in the proof of Proposition 5.3 of the $n$-tuple of linked quaternion algebras can also be taken to be the symbol $\alpha\frac{d\beta_{1}}{\beta_{1}}\wedge\dots\wedge\frac{d\beta_{n}}{\beta_{n}}$ in $H_{2}^{n+1}(F)$. One could hope that a similar phenomenon holds for odd primes, but unfortunately, the symbol $\alpha\frac{d\beta_{1}}{\beta_{1}}\wedge\dots\wedge\frac{d\beta_{n}}{\beta_{n}}$ in $H_{p}^{n+1}(F)$ is not an invariant of the sequence $([\alpha,\beta_{1})_{p,F},\dots,[\alpha,\beta_{n})_{p,F})\in\operatorname{LCA}_{p,n}(F)$ for $p>2$ as shown in [15, Remark 3.6].
@@ 6. The $\operatorname{Dec}_{p^{m},n}$ functor
In this section, we say a few words about the functor $\operatorname{Dec}_{p^{m},n}$ mapping each field $F$ containing $k$ to the subset of algebras in $\operatorname{Alg}_{p^{mn},p^{m}}(F)$ that are isomorphic to tensor products of $n$ cyclic algebras of degree $p^{m}$ over $F$. The essential dimension and $p$-dimension of this functor were studied in [36] for $m=1$, and it was concluded that they are both equal to $n+1$.
@@
There is a surjective map from $\operatorname{Dec}_{p^{m},n}(F)$ to $\operatorname{Dec}_{p,n}(F)$ given by mapping each $(\omega_{1},\dots,\omega_{m}),\beta)_{p^{m},F}$ to $\omega_{1},\beta)_{p,F}$. Hence $\operatorname{ed}_{p}(\operatorname{Dec}_{p^{m},n})\geqslant\operatorname{ed}_{p}(\operatorname{Dec}_{p,n})=n+1$ by [36]. On the other hand, by [43, Theorem 1], the essential $p$-dimension of $(\mathbb{Z}/p^{m}\mathbb{Z})^{\times n}$ over fields of characteristic $p$ is 1, which means that every $n$-tuple $(\tau_{1},\dots,\tau_{n})$ of classes in $H_{p^{m}}^{1}(F)$ descends to a field $L$ of transcendence degree 1 over $k$. Hence a tensor product of $n$ cyclic algebras $\tau_{1},\beta_{1})_{p^{m},F}\otimes\cdots\otimes[\tau_{n},\beta_{n})_{p^{m},F}$ of degree $p^{m}$ over a given $F$ descends to $L(\beta_{1},\ldots,\beta_{n})$, showing that the essential dimension is at most $n+1$. ∎
See Section 9 for more information on $\operatorname{Dec}_{p,n}$.
@@ 7. Cohomology Groups
In the good characteristic case, Galois Cohomology gives a description of the $n$-torsion part of the Brauer group; ${{}_{n}{\operatorname{Br}}}(F)\cong H^{2}(F,\mu_{n})$. In the bad characteristic case, the analogous cohomological construction is Kato-Milne cohomology for prime-torsion (see [30]) and Izhboldin groups in $p^{m}$-torsion (see [26] and [3]). As pointed out to us by J. P. Tignol, though they are called “Izhboldin groups”, they were previously known to Kato and used by him (see [29, Section 2, Corollary 4 to Proposition 2]). Both constructions utilize differential forms and are described here.
@@
The construction of Kato-Milne cohomology is as follows: we start with a field $F$ of $\operatorname{char}(F)=p$. Take $\Omega_{F}^{n}$ the ring of $n$-fold differential forms (i.e., sums of terms of the form $a_{1}db_{1}\wedge\dots\wedge db_{n}$ subject to the relations $d(bc)=bdc+cdb$ and $db\wedge db=0$), and consider the map $\wp:\Omega_{F}^{n}\rightarrow\Omega_{F}^{n}/d\Omega_{F}^{n-1}$ defined by $a\frac{db_{1}}{b_{1}}\wedge\dots\wedge\frac{db_{n}}{b_{n}}\mapsto(a^{p}-a)\frac{db_{1}}{b_{1}}\wedge\dots\wedge\frac{db_{n}}{b_{n}}$. The group $H_{p}^{n+1}(F)$ is the cokernel of this map, and it is the analogous group to $H^{n+1}(F,\mu_{p}^{\otimes n})$ in the more common case of fields $F$ of characteristic not $p$ containing primitive $p$th roots of unity (see [22, Section 101]). In particular, for $p=2$, $H_{2}^{n+1}(F)$ is isomorphic to $I_{q}^{n+1}(F)/I_{q}^{n+2}(F)$, where $I_{q}^{n+1}(F)$ is the subgroup of the Witt group of nonsingular quadratic forms over $F$ generated by generalized $(n+1)$-fold Pfister forms. The isomorphism is given by $a\frac{db_{1}}{b_{1}}\wedge\dots\wedge\frac{db_{n}}{b_{n}}\mapsto\langle\!\langle b_{1},\dots,b_{n},a]\!]$ (see also [22, Section 16]). For $n=1$, $H_{p}^{2}(F)$ is isomorphic to ${{}_{p}{\operatorname{Br}}}(F)$, and the isomorphism is given by $a\frac{db}{b}\mapsto[a,b)_{p,F}$ (see Section 8). The symbol length of a class in $H_{p}^{n+1}(F)$ is the number of terms of the form $a\frac{db_{1}}{b_{1}}\wedge\dots\wedge\frac{db_{n}}{b_{n}}$ needed to express it, and this number is always bounded from above by $\binom{r}{n}$ when $r$ is the $p$-rank of $F$ (see [20, Corollary 3.4]). Recall that when $F$ is a finitely generated field containing $k$ (and thus $\operatorname{char}(F)=p$), $F$ is a vector space over $F^{p}$ of dimension $p^{r}$ for some $r<\infty$, and this $r$ is called the “$p$-rank” of $F$, and that when $F$ is of transcendence degree $n$ over $k$, its $p$-rank is $n$ (see [11, Chapter V, Section 16.6, Corollary 3]). Knowing the symbol length may say something about the decomposability of the algebra, which is the trick applied in [20] to provide lower bounds for the essential dimension of $\operatorname{Alg}_{p^{n},p}$.
@@
The group $\nu(n)_{F}$, defined to be the kernel of $\wp$, is the one corresponding to Milnor $K$-theory (i.e., the group $K_{n}(F)/pK_{n}(F)$), and it is the group $H^{n}(F,\mu_{p}^{\otimes n})$ in the language of [22, Section 101]. For $p=2$, this group is isomorphic to $I^{n}F/I^{n+1}F$ where $I^{n}F$ is the group generated by generalized bilinear $n$-fold Pfister forms inside the additive Witt group $WF$. The Witt group is actually a ring, because the tensor product is defined, and more generally, there is a cup product that maps pairs $(A,B)\in\nu(\ell)_{F}\times\nu(m)_{F}$ to $A\wedge B\in\nu(\ell+m)_{F}$. Similarly, there is a cup product that maps pairs $(A,B)\in H_{p}^{\ell}(F)\times\nu(m)_{F}$ to $A\wedge B\in H_{p}^{\ell+m}(F)$. But note that since $H_{p}^{m}(F)$ and $\nu(m)_{F}$ need not be isomorphic, there is no cup product for pairs in $H_{p}^{\ell}(F)\times H_{p}^{m}(F)$.
In the good characteristic case, the cup product for pairs in $H^{\ell}(F,\mu_{p})\times H^{m}(F,\mu_{p})$ is defined, and it is applied (see [38, Example 3.7]) to produce the lower bound of $2n$ for the essential dimension of $\operatorname{Alg}_{p^{n},p}$: there is a well-defined map from tensor products $(\alpha_{1},\beta_{1})_{p,F}\otimes\dots\otimes(\alpha_{n},\beta_{n})_{p,F}$ of $n$ cyclic algebras of degree $p$ over a field $F$ of characteristic 0 containing primitive $p$th roots of unity to the cup product of these algebras $(\alpha_{1},\beta_{1})\cup\dots\cup(\alpha_{n},\beta_{n})=(\alpha_{1},\beta_{1},\dots,\alpha_{n},\beta_{n})\in H^{2n}(F,\mu_{p})$. Since the cohomological $p$-dimension of a field $E$ with $k\subseteq E\subseteq F$ is bounded from above by the transcendence degree of $E$ over $k$, the lower bound readily follows.
We cannot repeat the same argument in the bad characteristic case, because there is no cup product. However, the lower bound for the essential dimension of $\operatorname{Alg}_{p^{n},p}$ in the good case mentioned above is $2n$ which corresponds to the essential dimension of $\operatorname{Dec}_{p,n}$ in this case. Similarly, in [36], a lower bound of $n+1$ was produced for the essential dimension of $\operatorname{Alg}_{p^{n},p}$ in the bad case, which again coincides with the essential dimension of $\operatorname{Dec}_{p,n}$ in this case. See Proposition 9.2 for an explanation of this fact.
@@
The Izhboldin groups $H_{p^{m}}^{n+1}(F)$ can be obtained in a similar way as a quotient of $W_{m}\Omega^{n}_{F}$ ([3, Theorem 2.27]), which is the ring of truncated Witt vectors of length $m$ over $\Omega^{n}_{F}$ (see [27, Chapter 8, Section 10]), but a neater description in terms of generators and relations is the following ([26]): take the additive group $W_{m}(F)\otimes\underbrace{F^{\times}\otimes\dots\otimes F^{\times}}_{n\ \text{times}}$ modulo the relations
@@
The group $H_{p^{m}}^{1}(F)$ is of particular interest because it describes the degree $p^{m}$ cyclic extensions of $F$, and is in fact $W_{m}F/\wp(W_{m}F)$. The essential dimension of the functor $H_{p^{m}}^{1}$ is the same as that of the finite group $\mathbb{Z}/p^{m}\mathbb{Z}$ and is unknown in general, except for the cases of $m=1$ and $2$, for which it is 1 and 2 respectively. See [31] for further details, where it is also proven that $\operatorname{ed}(\mathbb{Z}/p^{m}\mathbb{Z})\leqslant m$ in general, and an equality between the two is conjectured. However, the essential $p$-dimension of the functor $H_{p^{m}}^{1}$ is known to be 1 (see [43]), which provides for an interesting case where the essential $p$-dimension is strictly smaller than the essential dimension.
@@
Given $(a,b)\in H^{1}_{p^{2}}(F)$, there exists a prime-to-$p$ extension $L$ of $F$ of degree at most $p^{2}-p+1$ such that $(a,b)_{L}=(c,0)$ for a suitable $c\in L$.
@@
The second slot is thus a monic polynomial of degree $p^{2}-p+1$, and therefore has a root in some prime-to-$p$ extension $L$ of $F$ of degree at most $p^{2}-p+1$. When $p=2$,
@@
Question 7.2.
Can the argument in Proposition 7.1 be extended to show that $\operatorname{ed}_{p}(H^{1}_{p^{m}})=1$ for $m>2$?